McGilla Golf has decided to sell a new line of golf clubs. The company would like to know the sensitivity of NPV to changes in the price of the new clubs and the quantity of new clubs sold. The clubs will sell for $795 per set and have a variable cost of $355 per set. The company has spent $200,000 for a marketing study that determined the company will sell 65,000 sets per year for seven years. The marketing study also determined that the company will lose sales of 11,000 sets of its high-priced clubs. The high-priced clubs sell at $1,165 and have variable costs of $625. The company will also increase sales of its cheap clubs by 13,000 sets. The cheap clubs sell for $385 and have variable costs of $175 per set. The fixed costs each year will be $10,050,000. The company has also spent $1,500,000 on research and development for the new clubs. The plant and equipment required will cost $37,800,000 and will be depreciated on a straight-line basis. The new clubs will also require an increase in net working capital of $2,200,000 that will be returned at the end of the project. The tax rate is 25 percent, and the cost of capital is 13 percent. What is the sensitivity of the NPV to each of these variables?
Answer #1

1.
Sensitivity wrt Price=65000*(1-25%)/13%*(1-1/1.13^7)=215602.258595

2.
Sensitivity wrt Quantity=(795-355)*(1-25%)/13%*(1-1/1.13^7)=1459.461443